# Exact holomorphic solution of the Hodge-dual minimal surface: a confining quantum QCD string

### v1

@@ Abstract
We present an exact analytic solution for the quaternionic Hodge–dual minimal surface in four dimensions, which provides a geometric fixed point of the Yang-Mills Wilson loop equation. The area of this surface is expressed through the eigenvalues of the Douglas–Gram matrix of the boundary loop, and is identical for the self–dual and anti–self–dual chiralities related by parity. We prove the Confinement Theorem by showing that the effective string tension $\sigma_{\mathrm{eff}}[C]$ (ratio of our Hodge dual area to a Euclidean one bounded by the same loop) is bounded as $\sqrt{2}\leq\sigma_{\mathrm{eff}}[C]\leq 2\sqrt{2}$, remaining finite and positive for all regular loops. Consequently, the Wilson loop decays exponentially with the Euclidean area, providing a purely geometric proof of confinement. Quantization of the quadratic holomorphic action yields a one-loop exact quantum string, with a universal $\tfrac{1}{2}\ln(2\pi)$ correction for each of the two chiral strings. The worldsheet theory is one-loop exact; the quantum corrections vanish after factoring out the self-dual rotational gauge orbit, leaving a purely spectral, gauge-invariant action. This construction defines a solvable quantum Hodge–dual string theory reproducing the QCD area law with explicit tension dependence on the loop shape.
@@ 1 Introduction
Since the pioneering work of Ken Wilson [1], understanding confinement in Yang–Mills theory remains one of the central challenges of modern physics. Conventional approaches based on the local action obscure the geometric structure of the Wilson loop, where confinement is expected to emerge as an area law. In this work, we use the loop–space calculus [2] to reformulate the problem in purely geometric terms, leading to its analytic solution. In [2], the general Euler-Lagrange (EL) equations for the SD/ASD surface were derived, and an exact solution was found for the class of planar loops. We construct a Hodge-dual minimal surface, such that the corresponding area law exactly solves the loop equation for the Yang–Mills gradient flow. Here, we generalize this by presenting a complete holomorphic solution for arbitrary regular loops. This holomorphic solution supersedes the planar case constructed in Ref. [2], extending it to arbitrary loops and simultaneously correcting a typographical error $Y^{A}_{\mu}=0$ in the proof of the planar solution in that paper.
@@ 2 Basic definitions and duality equations
$S_{\chi}[C;X]$ $=\min_{X\in\mathcal{M}_{\rm SD/ASD}}\int_{\mathbb{D}}\!\mathrm{d}^{2}z\,\sqrt{\tfrac{1}{2}\,\Sigma_{\mu\nu}\Sigma_{\mu\nu}},$ (2.1)
@@
With $(\ast\Sigma)_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\rho\sigma}\Sigma_{\rho\sigma}$, the SD/ASD parts are $\Sigma^{(\pm)}_{\mu\nu}=\tfrac{1}{2}(\Sigma_{\mu\nu}\pm\ast\Sigma_{\mu\nu})$, satisfying $\ast\Sigma^{(\pm)}_{\mu\nu}=\pm\Sigma^{(\pm)}_{\mu\nu}$. The Euler-Lagrange variational equations for this functional were written down in [2]. These nonlinear differential equations were solved in that paper only for a special case of planar boundary loop $C$. In this paper, we generalize that solution for an arbitrary loop.
Self-dual area derivative. There is an internal $O(4)$ gauge invariance of our area element in addition to a physical $O(4)$ invariance of the problem in $\mathbb{R}^{4}$. We use that extra degree of freedom to provide the self-duality (SD or ASD) of the area derivative [2]. The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
@@
The loop equation for the Yang–Mills gradient flow was shown in [2] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
@@
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [2] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents SD or ASD tensor. This would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
@@ 3 Geometric Definition of the Physical Area
For a loop that is a composite of two (or more) separated or self-intersecting sub-loops, $C=C_{1}\uplus C_{2}$, the problem changes. The E-L equations and the Dirichlet BCs are local. The problem of finding the minimal surface $X$ for the boundary $C$ therefore decouples into two independent minimal surface problems: $X_{1}$ on domain $\mathbb{D}_{1}$ bounded by $Q_{1}C_{1}$, and $X_{2}$ on domain $\mathbb{D}_{2}$ bounded by $Q_{2}C_{2}$. The total action for the solution is the sum $S[X]=S[X_{1}]+S[X_{2}]$.
@@
To find the true maximum over "all compatible isometries" (per our definition), we must maximize over this larger group. This defines the area for a composite loop as $S_{\mathrm{phys}}[C_{1}\uplus C_{2}]=\max_{R_{1}\in O(4)}S[X_{1}[R_{1}C_{1}]]+\max_{R_{2}\in O(4)}S[X_{2}[R_{2}C_{2}]]$.
@@ 4 A Self-Dual Holomorphic Ansatz
where the second line follows from $\Sigma_{z_{1}\bar{z}_{1}}=\Sigma_{z_{2}\bar{z}_{2}}=\mathrm{i}a(z)$ and the definition $\Sigma_{z_{a}\bar{z}_{b}}=\mathrm{i}\sum_{A}V^{A}_{z_{a}}\overline{V^{A}_{z_{b}}}$.
@@
$S_{\chi}=2\sqrt{2}\int_{\mathbb{D}}\big(|f_{1}^{\prime}|^{2}+|f_{2}^{\prime}|^{2}\big)\mathrm{d}^{2}z.$ (4.4)
@@
Admissible loops.
Throughout we assume the boundary loop $C:S^{1}\to\mathbb{R}^{4}$ is admissible in the sense that the spectral expansion converges in the Douglas–Gram matrix, so that the Dirichlet identity leading to (4.6) is valid, and that the Fourier/Hilbert-transform manipulations used in this paper are justified.
@@ 5 Invariant Area and the Spectral Law
Let $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$ be the eigenvalues of $G$. The maximum in (5.1) is evaluated by first applying the cyclic property of the trace, which gives $\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}((R^{T}PR)G)$. We can define a new projector $P^{\prime}=R^{T}PR$. Since $P$ is a fixed rank-2 projector, $P^{\prime}$ is also a rank-2 projector. The group $O(4)$ acts transitively on the space of all 2-planes, $\operatorname{Gr}(2,4)$. This means that by maximizing over all $R\in O(4)$, the projector $P^{\prime}$ sweeps through the entire set of all possible rank-2 projectors. Therefore, our maximization is mathematically identical to $2\sqrt{2}\max_{P^{\prime}\in\operatorname{Gr}(2,4)}\operatorname{tr}(P^{\prime}G)$. By the Ky Fan theorem [3], this maximum is precisely the sum of the two largest eigenvalues, $\lambda_{3}+\lambda_{4}$.
@@
This calculation confirms that our physical definition of the invariant area (5.1) leads directly to the $O(4)$-invariant spectral law. For planar loops ($\operatorname{rank}G=2$), $\lambda_{1}=\lambda_{2}=0$, and this correctly specializes to $S_{\mathrm{phys}}[\mathcal{C}]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})=2\sqrt{2}\,\mathcal{D}[C]$, reproducing the claim of Theorem 1 of [2]. The optimal projector $P_{\ast}=v_{3}v_{3}^{\top}+v_{4}v_{4}^{\top}$ (where $v_{i}$ are eigenvectors of $G$) is the one that achieves this maximum.
@@ 6 Self-Duality of the Area Derivative
We begin with the generic formula from [2] for the first variation of the Hodge-dual area
@@
where $\hat{Q}_{\mu\nu}$ is the constant, antisymmetric tensor $\hat{Q}=\text{diag}(i\sigma_{2},i\sigma_{2})$. In the second derivative, by the chain rule, we use the standard Volterra derivative [4] of the loop with respect to its tangent, $\delta C(x)/\delta C^{\prime}(y)=\Theta(x-y)$. This tensor $\hat{Q}$ is manifestly SD (as $\ast\hat{Q}_{12}=\hat{Q}_{34}$ and $\ast\hat{Q}_{34}=\hat{Q}_{12}$), which proves the area derivative for $S_{\chi}[C;P]$ is self-dual. By the covariance argument above, this property extends to $S_{\mathrm{phys}}$.

### v2

# Geometric Solution of the Loop Equation and Quark Confinement in Planar QCD

@@ Abstract
We present an exact analytic area law for planar QCD satisfying the MM loop equation for arbitrary loops. The solution is constructed using a quaternionic Hodge-dual minimal surface in a 16-dimensional auxiliary space. Unlike random surfaces, this construction defines a topological string theory where the action is proportional to the area of the conformal map of the 2D disk, possessing no bulk quantum fluctuations.
The Wilson loop factorizes into a confining dressing factor and the perturbative term: $W[C]=W_{pert}[C]e^{-\kappa S[C]}$. This corresponds to a non-perturbative gluon condensate vacuum supporting planar graph fluctuations. We prove the inequality $S[C]\geq\sqrt{2}A[C]$ (where $A[C]$ is the Euclidean minimal area), establishing a sufficient condition for confinement.
The area law is defined explicitly as $S[C]=2\sqrt{2}(\lambda_{3}+\lambda_{4})$ using the two largest eigenvalues of the Douglas-Gram matrix. We demonstrate that this functional is an exact additive zero mode of the loop diffusion operator, protected algebraically by the Bianchi identity. This ensures global validity for all regular loops ($C\in H^{1/2}(\mathbb{S}^{1},\mathbb{R}^{4})$).
Matching this confining factor with the gluon condensate via the QCD Operator Product Expansion (OPE) determines the physical string tension $\kappa\sim\Lambda_{QCD}^{2}$ (up to a universal numerical factor). This framework provides a first-principles derivation of confinement in large-$N_{c}$ QCD and new tools for the high-lying meson spectrum.
@@ 1 Introduction
Since the pioneering work of Ken Wilson [Wilson1974Confinement], understanding confinement in Yang–Mills theory remains one of the central challenges of modern physics. Conventional approaches based on the local action obscure the geometric structure of the Wilson loop, where confinement is expected to emerge as an area law. In this work, we use the loop–space calculus [migdal2025SQYMflow] to reformulate the problem in purely geometric terms, leading to its analytical solution. In [migdal2025SQYMflow], the general Euler-Lagrange (EL) equations for the SD/ASD surface were derived, and an exact solution was found for the class of planar loops. We construct a Hodge-dual minimal surface, such that the corresponding area law exactly solves the loop equation for the Yang–Mills gradient flow. Here, we generalize this by presenting a complete holomorphic solution for arbitrary regular loops (with Fourier coefficients decreasing faster than $1/n$). This holomorphic solution supersedes the planar case constructed in Ref. [migdal2025SQYMflow], extending it to arbitrary loops and simultaneously correcting a typographical error $Y^{A}_{\mu}=0$ in the proof of the planar solution in that paper.
@@ 2 Basic definitions and duality equations
$S_{\chi}[C;X]$ $=\min_{X\in\mathcal{M}_{\rm SD/ASD}}\int_{\mathcal{D}}\!\mathrm{d}^{2}z\,\sqrt{\tfrac{1}{2}\,\Sigma_{\mu\nu}\Sigma_{\mu\nu}},$ (2.1)
@@
With $(\ast\Sigma)_{\mu\nu}=\tfrac{1}{2}\epsilon_{\mu\nu\rho\sigma}\Sigma_{\rho\sigma}$, the SD/ASD parts are $\Sigma^{(\pm)}_{\mu\nu}=\tfrac{1}{2}(\Sigma_{\mu\nu}\pm\ast\Sigma_{\mu\nu})$, satisfying $\ast\Sigma^{(\pm)}_{\mu\nu}=\pm\Sigma^{(\pm)}_{\mu\nu}$. The Euler-Lagrange variational equations for this functional were written down in [migdal2025SQYMflow]. These nonlinear differential equations were solved in that paper only for a special case of planar boundary loop $C$. In this paper, we generalize that solution for an arbitrary loop.
Self-dual area derivative. There is an internal $O(4)$ gauge invariance of our area element in addition to a physical $O(4)$ invariance of the problem in $\mathbb{R}^{4}$. We use that extra degree of freedom to provide the self-duality (SD or ASD) of the area derivative [migdal2025SQYMflow]. The area derivative for a general functional $W[C]$ was defined in that paper as a discontinuity of the second variation of the area
@@
The loop equation for the Yang–Mills gradient flow was shown in [migdal2025SQYMflow] to reduce to the loop space diffusion equation, which allows the solution $W=\exp\left(-\kappa S[C]\right)$ provided that $S[C]$ is a zero mode of this equation
@@
The last equation (Bianchi identity) is valid at the kinematical level as a Jacobi identity for a triple commutator (see [migdal2025SQYMflow] for a proof). The matrix in the Dirichlet boundary conditions (BC) for $X\sim QC$ should be chosen so that the area derivative represents SD or ASD tensor. This would make the exponential of the Hodge-dual minimal area an exact solution to the loop equation.
@@ 3 Geometric Definition of the Physical Area
For a loop that is a composite of two (or more) separated or self-intersecting sub-loops, $C=C_{1}\uplus C_{2}$, the problem changes. The E-L equations and the Dirichlet BCs are local. The problem of finding the minimal surface $X$ for the boundary $C$ therefore decouples into two independent minimal surface problems: $X_{1}$ on domain $\mathcal{D}_{1}$ bounded by $Q_{1}C_{1}$, and $X_{2}$ on domain $\mathcal{D}_{2}$ bounded by $Q_{2}C_{2}$. The total action for the solution is the sum $S[X]=S[X_{1}]+S[X_{2}]$.
@@
To find the true maximum over "all compatible isometries" (per our definition), we must maximize over this larger group. This defines the area for a composite loop as $S_{\mathrm{phys}}[C_{1}\uplus C_{2}]=\max_{R_{1}\in O(4)}S[X_{1}[R_{1}C_{1}]]+\max_{R_{2}\in O(4)}S[X_{2}[R_{2}C_{2}]]=S_{\mathrm{phys}}[C_{1}]+S_{\mathrm{phys}}[C_{2}]$.
Finally, this maximized minimal projection area depends on the parametrization of the bounding curve, unlike the Wilson loop which is reparametrization invariant by definition. This last ambiguity is eliminated in the same way as in the Euclidean minimal surface: by minimizing the resulting functional over the group of orientation-preserving diffeomorphisms of the circle, $C(\theta)\to C(\phi(\theta))$ with $\phi^{\prime}>0$. Combining these steps, the full definition for a loop decomposable into sub-loops $C_{i}$ is:
$S[C]=\sum_{C_{i}\in C}$
$\min_{\phi\in\mathrm{Diff}^{+}(S^{1})}\max_{R\in O(4)}\min_{\begin{subarray}{c}X:\mathbb{D}_{i}\to\mathbb{R}^{4}\otimes\mathbb{R}^{4}\\ X|_{\partial\mathbb{D}_{i}}=QR(C_{i}\circ\phi)\end{subarray}}\int_{\mathbb{D}_{i}}\lVert dX^{A}\wedge dX^{A}\rVert$ (3.2)
@@ 4 A Self-Dual Holomorphic Ansatz
, where the second line follows from $\Sigma_{z_{1}\bar{z}_{1}}=\Sigma_{z_{2}\bar{z}_{2}}=\mathrm{i}a(z)$ and the definition $\Sigma_{z_{a}\bar{z}_{b}}=\mathrm{i}\sum_{A}V^{A}_{z_{a}}\overline{V^{A}_{z_{b}}}$.
@@
$S_{\chi}=2\sqrt{2}\int_{\mathcal{D}}\big(|f_{1}^{\prime}|^{2}+|f_{2}^{\prime}|^{2}\big)\mathrm{d}^{2}z.$ (4.4)
@@
Regular loops.
Throughout we assume the boundary loop $C:S^{1}\to\mathbb{R}^{4}$ is regular in the sense that the spectral expansion converges in the Douglas–Gram matrix, so that the Dirichlet identity leading to (4.6) is valid, and that the Fourier/Hilbert-transform manipulations used in this paper are justified. This corresponds to $C\in H^{\tfrac{1}{2}}(S^{1},\mathbb{R}^{4})$
@@ 5 Invariant Area and the Spectral Law
Let $0\leq\lambda_{1}\leq\cdots\leq\lambda_{4}$ be the eigenvalues of $G$. The maximum in (5.1) is evaluated by first applying the cyclic property of the trace, which gives $\max_{R\in O(4)}2\sqrt{2}\,\operatorname{tr}((R^{T}PR)G)$. We can define a new projector $P^{\prime}=R^{T}PR$. Since $P$ is a fixed rank-2 projector, $P^{\prime}$ is also a rank-2 projector. The group $O(4)$ acts transitively on the space of all 2-planes, $\operatorname{Gr}(2,4)$. This means that by maximizing over all $R\in O(4)$, the projector $P^{\prime}$ sweeps through the entire set of all possible rank-2 projectors. Therefore, our maximization is mathematically identical to $2\sqrt{2}\max_{P^{\prime}\in\operatorname{Gr}(2,4)}\operatorname{tr}(P^{\prime}G)$. By the Ky Fan theorem [Horn:2012:MatrixAnalysis:2nd], this maximum is precisely the sum of the two largest eigenvalues, $\lambda_{3}+\lambda_{4}$.
@@
This calculation confirms that our physical definition of the invariant area (5.1) leads directly to the $O(4)$-invariant spectral law. For planar loops ($\operatorname{rank}G=2$), $\lambda_{1}=\lambda_{2}=0$, and this correctly specializes to $S_{\mathrm{phys}}[\mathcal{C}]=2\sqrt{2}\,(\lambda_{3}+\lambda_{4})=2\sqrt{2}\,\mathcal{D}[C]$, reproducing the claim of Theorem 1 of [migdal2025SQYMflow]. The optimal projector $P_{\ast}=v_{3}v_{3}^{\top}+v_{4}v_{4}^{\top}$ (where $v_{i}$ are eigenvectors of $G$) is the one that achieves this maximum.
5.1 Projector Alignment
The physical area $S_{phys}$ is realized by maximizing the base functional $S_{\chi}$ over the global $\mathrm{O}(4)$ rotation group $\mathbf{R}$: $S_{phys}=\max_{\mathbf{R}\in\mathrm{O}(4)}2\sqrt{2}\operatorname{tr}(\mathbf{P}(\mathbf{RGR}^{\top}))$. The initial fixed projector $\mathbf{P}$ is defined by the orthonormal basis vectors $\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ associated with the Hodge parity $\chi=\pm 1$:
$\mathbf{u}_{1}$ $=\frac{1}{\sqrt{2}}(\mathbf{e}_{1}+\mathbf{e}_{2})$ (5.3)
$\mathbf{u}_{2}$ $=\frac{1}{\sqrt{2}}(\mathbf{e}_{3}+\chi\mathbf{e}_{4})$ (5.4)
The optimal rotation $\mathbf{R}_{max}=\mathbf{R}_{2}\mathbf{R}_{1}$ maps this initial SD basis onto the optimal basis $\{\mathbf{v}_{3},\mathbf{v}_{4}\}$ corresponding to the largest eigenvalues of $\mathbf{G}$, $\lambda_{3}$ and $\lambda_{4}$.
5.1.1 1. Projection and Orthogonalization (the combined operation $\mathbf{R}_{1}$
The first operation $\mathbf{R}_{1}$ starts with alignement of the initial plane $\operatorname{span}(\mathbf{u}_{1},\mathbf{u}_{2})$ with the optimal eigenspace $\mathcal{V}_{34}=\operatorname{span}(\mathbf{v}_{3},\mathbf{v}_{4})$. This is achieved by projecting $\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ orthogonal to the null space spanned by the two lowest eigenvectors $\{\mathbf{v}_{1},\mathbf{v}_{2}\}$. Let $\mathbf{P}_{12}=\mathbf{v}_{1}\mathbf{v}_{1}^{\top}+\mathbf{v}_{2}\mathbf{v}_{2}^{\top}$ be the projector onto this null space. The intermediate vectors $\mathbf{u}^{\prime}_{i}$ are given by: $\mathbf{u}^{\prime}_{i}=\operatorname{Normalize}\left[\mathbf{u}_{i}-\mathbf{P}_{12}\cdot\mathbf{u}_{i}\right]$
The second part of this operation $\mathbf{R}_{1}$ orthogonalizes these two vectors: $\mathbf{u}^{\prime\prime}_{2}=\operatorname{Normalize}\left[\mathbf{u}_{2}-(\mathbf{u}_{2}\cdot\mathbf{u}_{1})\mathbf{u}_{1}\right]$ After these two transformations, the basis vectors $\mathbf{u}_{1},\mathbf{u}_{2}$ become orthogonal basis in the Eigenspace ($\mathbf{R}_{2}$. Therefore, this combined transformation is an $O(4)$ rotation.
5.1.2 2. Alignment within the Eigenspace ($\mathbf{R}_{2}$)
The second rotation $\mathbf{R}_{2}$ performs an $\mathrm{O}(2)$ transformation within the optimal plane $\mathcal{V}_{34}$ to fix the alignment of the resulting basis. This step aligns the first intermediate vector $\mathbf{u}^{\prime}_{1}$ with the highest eigenvector $\mathbf{v}_{3}$: $\mathbf{R}_{2}:\quad\text{An }\mathrm{O}(2)\text{ rotation in }\mathcal{V}_{34}\text{ such that }\mathbf{R}_{2}\mathbf{u}^{\prime}_{1}\propto\mathbf{v}_{3}$ The combined rotation $\mathbf{R}_{max}$ defines the required boundary condition $\mathbf{X}=\mathbf{Q}\cdot\mathbf{R}_{max}\mathbf{C}$ for the minimal surface, which yields the spectral area $S_{phys}$.
@@ 6 Self-Duality of the Area Derivative
We begin with the generic formula from [migdal2025SQYMflow] for the first variation of the Hodge-dual area
@@
where $\hat{Q}_{\mu\nu}$ is the constant, antisymmetric tensor $\hat{Q}=\text{diag}(i\sigma_{2},i\sigma_{2})$. In the second derivative, by the chain rule, we use the standard Volterra derivative [Gelfand:1963:CalculusVariations] of the loop with respect to its tangent, $\delta C(x)/\delta C^{\prime}(y)=\Theta(x-y)$. This tensor $\hat{Q}$ is manifestly self-dual (as $\ast\hat{Q}_{12}=\hat{Q}_{34}$ and $\ast\hat{Q}_{34}=\hat{Q}_{12}$), which proves that the area derivative for $S_{\chi}[C;P]$ is self-dual. By the covariance argument above, this property extends to $S_{\mathrm{phys}}$.
@@
7 The Hodge parity flip
How does our solution change for the negative Hodge parity $\chi=-1$ (ASD instead of SD)? There are many equivalent forms of the ASD solutions, all equivalent after the isometry transformation $O(4)$ rotation of the matrix $Q\Rightarrow QR$ in the Dirichlet BC $X=QC$. The conical transformation is a reflection $X^{A}_{4}\Rightarrow-X^{A}_{4}$ which in complex basis corresponds to $z_{2}\leftrightarrow\bar{z}_{2}$ in our holomorphic ansatz (4.3). After reflection, the only nonzero components will be
$U^{4}_{z_{1}}=f_{1}(z),\hskip 18.49988ptU^{3}_{\bar{z}_{2}}=f_{1}(z),$
$U^{2}_{z_{1}}=f_{2}(z),\hskip 18.49988ptU^{1}_{\bar{z}_{2}}=f_{2}(z).$ (7.1)
As a result, the $(2,0)$ and $(0,2)$ components of the form $\Sigma$ will still be zero, and the $(0,1),(1,0)$ components change. Now the $\Sigma$ form will become $\Sigma=a(z)\tilde{B}$, where $\tilde{B}~e_{1}\wedge e_{2}-e_{4}\wedge e_{4}$ which has negative Hodge duality. The energy density $a(z)$ stays the same, and so does the area functional. The BC (4.5) change in an obvious way
$\Re f_{2}(e^{\mathrm{i}\theta})$ $=\frac{1}{\sqrt{2}}\big(C_{3}(\theta)-C_{4}(\theta)\big).$ (7.2)
which leaves the projection operator $P$ unchanged.
The area derivative in (6.4) will change in an obvious way, following the change of the area element $B\Rightarrow\tilde{B}$
$\delta S\propto\iint d\theta_{1}d\theta_{2}H(\theta_{1}-\theta_{2})\tilde{Q}(C(\theta_{1}),\delta C^{\prime}(\theta_{2}));$ (7.3)
$\tilde{Q}(A,B)=A_{2}B_{1}-A_{1}B_{2}-A_{3}B_{4}+A_{4}B_{3}$ (7.4)
As expected, the area does not depend on the Hodge parity, but the area derivative switches from SD to ASD under Hodge parity flip. The loop equation will still hold, the same way as for the SD solution.